A recent broadcast of a television show had a

1010

share, meaning that among

60006000

monitored households with TV sets in use,

1010 %

of them were tuned to this program. Use a 0.01 significance level to test the claim of an advertiser that among the households with TV sets in use, less than

1515 %

were tuned into the program. Identify the null hypothesis, alternative hypothesis, test statistic, P-value, conclusion about the null hypothesis, and final conclusion that addresses the original claim. Use the P-value method. Use the normal distribution as an approximation of the binomial distribution.

Identify the null and alternative hypotheses. Choose the correct answer below.

A.

H0 :

pequals=0.150.15

H1 :

pless than<0.150.15

B.

H0 :

pequals=0.850.85

H1 :

pnot equals≠0.850.85

C.

H0 :

pequals=0.150.15

H1 :

pgreater than>0.150.15

D.

H0 :

pequals=0.850.85

H1 :

pless than<0.850.85

E.

H0 :

pequals=0.850.85

H1 :

pgreater than>0.850.85

F.

H0 :

pequals=0.150.15

H1 :

pnot equals≠0.150.15

The test statistic is

zequals=nothing.

(Round to two decimal places as needed.)The P-value is

nothing.

(Round to four decimal places as needed.)

Identify the conclusion about the null hypothesis and the final conclusion that addresses the original claim.

Fail to reject

Reject

H0.

There

is

is not

sufficient evidence to support the claim that less than

1515 %

of the TV sets in use were tuned to the program.
Answer #1

A. H0: p=0.15, H1: p<0.15

We have p_ hat= sample proportion= 0.1

n= sample size= 6000

The test statistic, z= (p_ hat-p)/se

se=√p(1-p)/n= 0.0046

z= (0.1-0.15)/0.0046 = -10.85

p-value= P(z<-10.85) = 0.0000

As p-value is less than alpha= 0.01 so Reject H0.

There is sufficient evidence to support the claim that less than 15% of the TV sets in use were tuned to the program.